How many gallons of a 4% acid solution should be mixed with 20 gallons of a 10% acid solution to obtain a mixture that is 5% acid?
Answer by scott8148(6628)

let x="gal of 4%" __ x(4%)+20(10%)=(x+20)(5%) __ multiply by 100 to clear percentages

4x+200=5x+100 __ x=100